OpenStudy (anonymous):

find which ordered pairs from a given set are a part of the solution set for the inequality 2y+3x≤9 (-1,8) (0,0) (2,2) (3,2)

OpenStudy (anonymous):

You only need substitute the x and y values for the ones given in each ordered pair; if the inequality holds true then that ordered pair is a part of the solution set. (Or, you could solve the equation and see which values correspond to the ordered pairs given.)

OpenStudy (anonymous):

plug every point into the inequality 1. -3+16=13>9 no 2. 0+0=0 <9 yes 3. 2*2+3*2= 10>9 no 4. 3*3+2*2=13>9 no so only (0,0) satisfied

OpenStudy (anonymous):

thanks cong guo

OpenStudy (anonymous):

i always just need the first equation and i can go from there lol

OpenStudy (anonymous):

sometimes cant understand whats i am to do lol